OpenStudy (shadyboy):

4d + 7 = -17 A d= -6 B d= -2.5 C d= 2.5 D d= -4.25

OpenStudy (anonymous):

solve for d 4d + 7 = -17 -7 -7 4d = -24 divide by 4 to get d alone d= -6, so A hope this helps!

OpenStudy (shadyboy):

Thanks

OpenStudy (anonymous):

Sure (: